# Magnon topology driven by altermagnetism

### v1

@@ VII Edge states
Figure 7: (a) Magnon spectrum of our model on a strip with $N_{y}=250$ shells in the $\bm{\hat{y}}$ direction obtained for the DMI constant $D=0.4J$, AM coupling $J_{1}=0.1J$, and Zeeman field $H=H_{s}/2$. Panel (b) shows a snapshot of the reconstructed dynamics of the magnetic moments of the edge state, corresponding to point A (for details, see Appendix D and SM video [69]). Panels (c) and (d) show the distribution of wavefunction intensity in Eq. (23) across the shells for the points A,B, and C, respectively.
@@
across the strip. Here, functions $\psi_{q_{x},n_{y},\nu}$ and $\psi_{-q_{x},n_{y},\nu}$ encode the amplitude and polarization of the magnetic moments precession, see Appendix D for details. To compare the intensity distributions of the wavefunctions $P_{n_{y}}$ over the strip for different points (A, B, C) we utilize the normalization rule $\sum_{n_{y}=1}^{N_{y}}P_{n_{y}}=1$ for all cases. As shown in Figs. 7(c) and 7(d), representative in-gap states [points A and B in Fig. 7(a)] are localized at the different edges of the strip, while a bulk state [e.g., point C in Fig. 7(a)] exhibits almost uniform distribution across the full width, with the average intensity $\overline{P}_{n_{y}}=1/N_{y}$. Additionally, using Eq. (B) we can reconstruct the real-space dynamics of the magnetic moments driven by Landau-Lifshitz equations for the edge and bulk states marked in Fig. 7(a) by points A and C, respectively (see SM videos  [^2] In supplemental materials we provide two movies which illustrate dynamics of magnetic moments for points A and C shown in Fig. 7(a). [The link will be provided by the editor.]).
@@ Acknowledgements.
We thank Fakher F. Assaad, Pedro M. Cônsoli, Alex Hickey, Oleg Janson, Toshihiro Sato for fruitful discussions. S. K., V. K., and J. v. d. B. acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project-ids 390858490 and 392019). J. v. d. B. also acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the Collaborative Research Center “Correlated Magnetism: From Frustration to Topology” (SFB 1143, project-id 247310070). K. Y. acknowledges financial support by the Deutsche Forschungsgemeinschaft (DFG, project-id YE 232/2-1).
@@ Appendix A Quantum linear spin-wave Hamiltonian and its diagonalization
$\bm{A}_{\bm{q}},\bm{B}_{\bm{q}}$ are $2\times 2$ matrices with entries $A^{\alpha\beta}_{\bm{q}},B^{\alpha\beta}_{\bm{q}}$ respectively. Note that $\bm{M}_{\bm{q}}$ is a Hermitian matrix as ${\cal H}_{2}$ is Hermitian. The additional trace term, $-\frac{1}{2}\sum_{\bm{q}}{\rm Tr}{\bm{A}}_{\bm{q}}$, arising from the bosonic commutation relations, is a $O(S)$ constant quantum contribution to the classical ground state energy $E_{\rm cl}$.
@@ A.1.1 Step-I
where $r_{1}=r_{6}=\sqrt{\frac{A^{11}_{\bm{q}}+Q_{2}}{Q_{1}+Q_{2}}}$, $r_{2}=\bar{r}_{5}=\sqrt{\frac{A^{11}_{\bm{q}}-Q_{1}}{Q_{1}+Q_{2}}}\,\frac{A^{12}}{\left|A^{12}\right|}$, $r_{3}=\bar{r}_{8}=-r_{1}\,\frac{A^{21}}{\left|A^{12}\right|}$, $r_{4}=r_{7}=-|r_{2}|$. This transformation satisfies $\bm{T}_{I}^{\dagger}\bm{\eta}\bm{T}_{I}=\bm{\eta}$ and transforms $\bm{M}_{\bm{q}}$ to
@@ Appendix B Spin-wave dynamics in the framework of the classical Landau-Lifshitz formalism
which coincides with the eiegnvalue problem in Eq. (76). The eigenvalues $(E_{+},E_{-},-E_{+},-E_{-})$ and the corresponding eigenvectors are discussed in sections A.1, and A.2. If the spin wave with wave-vector $\bm{q}$ is excited, then the real-space dynamics of magnetic moments in each of the sublattices is as follows
@@ Appendix C Altermagnetic limit
Figure 8: (a) Distribution of the amplitude of the magnon magnetic moment along the isolines of constant energy $E_{\bm{q}}^{\pm}=\text{const}$ in the limit of the Zeeman field $H=0$ and for the AM coupling $J_{1}=-0.04J$. Panels (b) and (c) show the distribution of magnon magnetic moment (106) within the Brillouin zone plotted. Panels (b) and (c) correspond to the upper and lower branches, respectively.
@@ Appendix D Edge states in the framework of the classical Landau-Lifshitz formalism
Magnetic Hamiltonian of the strip $\mathcal{H}^{\mathrm{str}}=\mathcal{H}^{\mathrm{str}}_{\textsc{afm}}+\mathcal{H}^{\mathrm{str}}_{\textsc{alt}}+\mathcal{H}^{\mathrm{str}}_{\textsc{dmi}}+\mathcal{H}^{\mathrm{str}}_{\textsc{z}}$ comprises the antiferroamgnetic
@@
$\begin{split}\bm{S}_{1}=&E_{0}\begin{bmatrix}1-\frac{J_{1}}{4J}&-\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&0&\mathscr{C}_{q_{x}}\\ -\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&\mathscr{F}-\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}&\mathscr{C}_{q_{x}}&0\\ 0&-\mathscr{C}_{q_{x}}&-1+\frac{J_{1}}{4J}&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}\\ -\mathscr{C}_{q_{x}}&0&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&-\mathscr{F}+\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}\end{bmatrix},\quad\bm{R}=E_{0}\begin{bmatrix}\frac{J_{1}}{4J}&-\mathscr{A}_{q_{x}}-\mathscr{B}_{q_{x}}&0&\mathscr{C}_{q_{x}}\\ 0&0&0&0\\ 0&-\mathscr{C}_{q_{x}}&-\frac{J_{1}}{4J}&\mathscr{A}_{q_{x}}-\mathscr{B}_{q_{x}}\\ 0&0&0&0\end{bmatrix},\\ \bm{S}=&E_{0}\begin{bmatrix}1-\frac{J_{1}}{2J}&-\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&0&\mathscr{C}_{q_{x}}\\ -\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&1-\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}&\mathscr{C}_{q_{x}}&0\\ 0&-\mathscr{C}_{q_{x}}&-1+\frac{J_{1}}{2J}&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}\\ -\mathscr{C}_{q_{x}}&0&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&-1+\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}\end{bmatrix},\quad\bm{L}=E_{0}\begin{bmatrix}\frac{J_{1}}{4J}&0&0&0\\ -\mathscr{A}_{q_{x}}-\mathscr{B}_{q_{x}}&0&\mathscr{C}_{q_{x}}&0\\ 0&0&-\frac{J_{1}}{4J}&0\\ -\mathscr{C}_{q_{x}}&0&\mathscr{A}_{q_{x}}-\mathscr{B}_{q_{x}}&0\end{bmatrix},\\ \bm{S}_{2}=&E_{0}\begin{bmatrix}\mathscr{F}-\frac{J_{1}}{4J}&-\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&0&\mathscr{C}_{q_{x}}\\ -\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&1-\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}&\mathscr{C}_{q_{x}}&0\\ 0&-\mathscr{C}_{q_{x}}&-\mathscr{F}+\frac{J_{1}}{4J}&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}\\ -\mathscr{C}_{q_{x}}&0&\mathscr{A}_{q_{x}}+\mathscr{B}_{q_{x}}&-1+\frac{J_{1}}{J}\sin^{2}\frac{q_{x}a_{0}}{2}\end{bmatrix}.\end{split}$ (113)

### v2

@@ VII Edge states
Figure 7: (a) Magnon spectrum of our model on a strip [by diagonalizing $\bm{H}_{q_{x}}^{\rm str}$ in Eq. (112)] with $N_{y}=250$ shells in the $\bm{\hat{y}}$ direction obtained for the DMI constant $D=0.4J$, AM coupling $J_{1}=0.1J$, and Zeeman field $H=H_{s}/2$. Panel (b) shows a snapshot of the reconstructed dynamics of the magnetic moments of the edge state, corresponding to point A (for details, see Appendix D and SM video [70]). Panels (c) and (d) show the distribution of wavefunction intensity in Eq. (23) across the shells for the points A,B, and C, respectively.
@@
across the strip. Here, functions $\psi_{q_{x},n_{y},\nu}$ and $\psi_{-q_{x},n_{y},\nu}$ encode the amplitude and polarization of the magnetic moments precession, see Appendix D for details. To compare the intensity distributions of the wavefunctions $P_{n_{y}}$ over the strip for different points (A, B, C) we utilize the normalization rule $\sum_{n_{y}=1}^{N_{y}}P_{n_{y}}=1$ for all cases. As shown in Figs. 7(c) and 7(d), representative in-gap states [points A and B in Fig. 7(a)] are localized at the different edges of the strip, while a bulk state [e.g., point C in Fig. 7(a)] exhibits almost uniform distribution across the full width, with the average intensity $\overline{P}_{n_{y}}=1/N_{y}$. Additionally, using Eq. (B) we can reconstruct the real-space dynamics of the magnetic moments driven by Landau-Lifshitz equations for the edge and bulk states marked in Fig. 7(a) by points A and C, respectively (see SM videos [70]).
@@ VIII Conclusions
We have focused here on a minimal model where the next-nearest-neighbor couplings act along a single direction per sublattice, i.e., a nonzero coupling $J_{1}$ within the checkerboard plaquettes and none in the empty ones. One can, however, also include an additional coupling $J_{2}$ along the diagonals of the empty plaquettes. As long as $J_{2}\neq J_{1}$, the system continues to exhibit altermagnetism, i.e., in the limit of $H=D=0$, the two sublattices can still be mapped to one another by $C_{4}$ lattice-rotation about the $\bm{\hat{z}}$ axis passing through either a $J_{1}$ or $J_{2}$ bond center combined with time-reversal flipping the spins. Essentially, the model has same symmetries for $J_{2}=0$ case and the case of nonzero $J_{2}\neq J_{1}$. We find that the model still hosts topological magnons with Chern numbers $|C^{\pm}|=1$, nonzero thermal Hall response, and chiral edge modes in a finite strip geometry. The gap between the magnon branches at X or Y now depends on the difference between the two couplings ($|J_{1}-J_{2}|$), which quantifies the strength of altermagnetism. While the thermal Hall conductivity may change quantitatively from the $J_{2}=0$ case, all qualitative physics of topology and transport remain unchanged.
@@ Acknowledgements.
We thank F. F. Assaad, P. M. Cônsoli, A. Hickey, O. Janson, T. Sato for fruitful discussions. S. K., V. K., and J. v. d. B. acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project-ids 390858490 and 392019). J. v. d. B. also acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the Collaborative Research Center “Correlated Magnetism: From Frustration to Topology” (SFB 1143, project-id 247310070). K. Y. acknowledges financial support by the Deutsche Forschungsgemeinschaft (DFG, project-id YE 232/2-1).
@@ Appendix A Quantum linear spin-wave Hamiltonian and its diagonalization
$\bm{A}_{\bm{q}},\bm{B}_{\bm{q}}$ are $2\times 2$ matrices with entries $A^{\alpha\beta}_{\bm{q}},B^{\alpha\beta}_{\bm{q}}$ respectively. Note that $\bm{M}_{\bm{q}}$ is a Hermitian matrix as ${\cal H}_{2}$ is Hermitian. The additional trace term, $\frac{1}{2}\sum_{\bm{q}}{\rm Tr}{\bm{A}}_{\bm{q}}$, arising from the bosonic commutation relations, is a $O(S)$ constant quantum contribution to the classical ground state energy $E_{\rm cl}$.
@@ A.1.1 Step-I
where $r_{1}=r_{6}=\sqrt{\frac{A^{11}_{\bm{q}}+Q_{2}}{Q_{1}+Q_{2}}}$, $r_{2}=\bar{r}_{5}=\sqrt{\frac{A^{11}_{\bm{q}}-Q_{1}}{Q_{1}+Q_{2}}}\,\frac{A^{12}_{\bm{q}}}{\left|A^{12}_{\bm{q}}\right|}$, $r_{3}=\bar{r}_{8}=-r_{1}\,\frac{A^{21}_{\bm{q}}}{\left|A^{12}_{\bm{q}}\right|}$, $r_{4}=r_{7}=-|r_{2}|$. This transformation satisfies $\bm{T}_{I}^{\dagger}\bm{\eta}\bm{T}_{I}=\bm{\eta}$ and transforms $\bm{M}_{\bm{q}}$ to
@@ Appendix B Spin-wave dynamics in the framework of the classical Landau-Lifshitz formalism
which coincides with the eigenvalue problem in Eq. (76). The eigenvalues $(E_{+},E_{-},-E_{+},-E_{-})$ and the corresponding eigenvectors are discussed in sections A.1, and A.2. If the spin wave with wave-vector $\bm{q}$ is excited, then the real-space dynamics of spins in each of the sublattices is as follows
@@ Appendix C Altermagnetic limit
Figure 8: (a) Distribution of the amplitude of the magnon magnetic moment (106) along the isolines of constant energy $E_{\bm{q}}^{\pm}=\text{const}$ in the limit of the Zeeman field $H=0$ and for the AM coupling $J_{1}=-0.04J$. Panels (b) and (c) show the distribution of magnon magnetic moment (106) within the Brillouin zone plotted. Panels (b) and (c) correspond to the upper and lower branches, respectively.
@@ Appendix D Edge states in the framework of the classical Landau-Lifshitz formalism
Magnetic Hamiltonian of the strip $\mathcal{H}^{\mathrm{str}}=\mathcal{H}^{\mathrm{str}}_{\textsc{afm}}+\mathcal{H}^{\mathrm{str}}_{\textsc{alt}}+\mathcal{H}^{\mathrm{str}}_{\textsc{dmi}}+\mathcal{H}^{\mathrm{str}}_{\textsc{z}}$ comprises the antiferromagnetic
@@
[…]